OpenStudy (jazzyfa30):

simplify \(\Large \dfrac{4x^3-2x^2+8x+8}{2x+1}\)

OpenStudy (anonymous):

well you can see that you have a common denominator so you can devide 2 on the top. Ex. \[2x^3 - x^2 + 4x + 4/x+ 1\] then Ex.2 \[(2x^3 - x^2) + (4x + 4) = x^2(2x - 1) + 4(x + 1)\] so you get left with \[(x^2 + 4)(2x - 1)\] because (x + 1) cancels out

OpenStudy (jazzyfa30):

ok

OpenStudy (anonymous):

are you ok

OpenStudy (jazzyfa30):

yes im fine y u ask

OpenStudy (anonymous):

just making sure

OpenStudy (jazzyfa30):

ok so do we divide all the x values by the constant and the variables

OpenStudy (amistre64):

if we run a synthetic division, x = -1/2 4x^3 -2x^2 +8x +8 0 -2 -2 -3 ----------------------- -1/2 ) 4 -4 6 R=5

OpenStudy (anonymous):

FriedRice, your solution is not correct. There is no common factor between the numerator and denominator. While you can factor 2 out of the numerator, factoring 2 out of the denominator would yield a denominator of \[x+\frac{ 1 }{ 2 }\]

OpenStudy (amistre64):

4x^2 -4x+6 + 5/(2x+1)

OpenStudy (anonymous):

If you perform long division, you end up with \[2x^2-2x+5+\frac{ 3 }{ 2x+1 }\]. As you can see, there is a remainder.

OpenStudy (jazzyfa30):

ok

OpenStudy (amistre64):

lol, yeah, mine a little off .... -4*-1/2 = 2, not -2

OpenStudy (amistre64):

4x^3 -2x^2 +8x +8 0 -2 2 -5 ----------------------- -1/2 ) 4 -4 10 R=3 thats better

OpenStudy (anonymous):

What course is this for, jazzyfa30? It would help for us to know what context the instruction "simply" is being used in.

OpenStudy (jazzyfa30):

algebra 2

OpenStudy (anonymous):

Are you working on polynomial long division, or do you believe this was supposed to be a neatly factored solution? As you can see, there is no "neat" solution to this beyond the long division I performed above.

OpenStudy (jazzyfa30):

polynomial

OpenStudy (amistre64):

if you flip the polys you can create a power series representation :)

OpenStudy (anonymous):

If long division of polynomials is where you are at in your course, then the quotient + remainder above would be the solution. If you were expected to factor and then simplify by canceling factors, this won't work in this case, as the polynomial in the numerator cannot be factored over integers. Thus, in that case, I would say 'this can't be simplified.'

OpenStudy (anonymous):

lol @ power series representation -- that'd be fun, right?

OpenStudy (amistre64):

8 -8x +14x^2-24x^3+48x^4 - .... --------------------- 1+2x ) 8 +8x -2x^2 +4x^3 -8-16x ------- -8x -2x^2 8x +16x^2 ----------- 14x^2 +4x^3 -14x^2 -28x^3 --------------- -24x^3 24x^3 +48x^4 -------------- then it evens out

OpenStudy (jazzyfa30):

thank u for stepping in and ir was a r=3 thingy at the top i wish i could the both of you a medal i am going to open 2 more questions and give you guys a medal Thank you @amistre and @ssilvestro

OpenStudy (jazzyfa30):

^ @amistre64

OpenStudy (amistre64):

youre welcome ;)

OpenStudy (anonymous):

Thank you!